OpenStudy (anonymous):

Find the number of years it will take for $27,000 to grow to $34262.61 at 6% interest compounded quarterly.

OpenStudy (kropot72):

The following equation is the starting point: \[34262.61=27000(1+\frac{0.06}{4})^{x}=27000\times (1.015)^{x}\] Can yhou see what the next step should be?

OpenStudy (anonymous):

So... 34262.61=(27405)^4x ??

OpenStudy (campbell_st):

no simplify the 1st equation \[34262.61 = 27000(1.015)^x\] then \[\frac{34262.61}{27000} = 1.015^x\] simplify the above equation do you know how to use logarithms...?

OpenStudy (anonymous):

It's one of those things if I'm pointed in the right direction I rememebr the way :S

OpenStudy (anonymous):

So am I just dividing both sides by 27000? And then logging it out?

OpenStudy (campbell_st):

yes... and with log laws \[\log(b^a) = alog(b)\] you need the above law for the right hand side... doesn't matter if its base 10 or base e logs and simplify the left hand side log (1.26899) = xlog(1.015) isolate x 1st then evaluate for the answer

OpenStudy (campbell_st):

when you find x.... it will be the number of quarters.... since the investment was compounded quarterly... divide by 4 for the years...

OpenStudy (anonymous):

4?

OpenStudy (campbell_st):

yes... I got 16 quarters or 4 years

OpenStudy (anonymous):

Yeah, x=16 So, I'm dividing that by four because its quarterly

OpenStudy (campbell_st):

thats correct.... hope it helps

OpenStudy (anonymous):

It does...just need to keep doing them. The logs have me hung up. Thank you!